Grade 11Organic Chemistry

19.5g of CH2FCOOH is dissolved in 500g of water.the depression in the freezing point of water observed is 1.0degree C.Ccalculate the vant hoff factor and dissociation constant of fluoroacitic acid??

malothu sushma
9 Years agoGrade 11

1 Answer

NETHAVATH ARCHANA
9 Years ago
HIII SUSHMA..Kf=1.86K kg/mol
tf=1c
w h2o=500g
Mh2o=18g/mol
w solute=19.5g
m solute=78g/mol
n=19.5/78=0.25mol
molality=0.25*1000/500=0.5
tf=i.kf.m
1=i.1.86.0.5
i=1/0.93
=1.075
alpha=1.075-1/1
=0.0752
ka=c.alpha2   
=0.5*0.752*0.752=2.8*10-3    ALL THE BEST